decompose into a partial fraction
3x^2+15x+8/(x+1)(x+2)(x-3)

Answer by stanbon(75887)

decompose into a partial fraction
3x^2+15x+8/(x+1)(x+2)(x-3)
---------------
3x^2+15x+8/(x+1)(x+2)(x-3) = A/(x+1) + B/(x+2) + C/(x-3)
= A(x+2)(x-3) + B(x+1)(x-3) + C(x+1)(x+2 / ......
= A(x^2-x-6) + B(x^2-2x-3) + C(x^2+3x+2)/ ....
------------------
A+B+C = 3
-A-2B+3C = 15
-6A-3B+2C = 8
-------------------
Using the Matrix function of a TI-83 I get:
A = 1
B = -2
C = 4
===========
Cheers,
Stan H.